Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
gosselin, chiummo, spinicelli, ruggi, bersanetti, melo - 23:10 Tuesday 26 April 2022 (55627) Print this report
Recentering of PR

This afternoon we started the activity about the PR centering.
We gave a first try by changing the setpoint of the BPC but it had not so much effect on the vertical shift of PR.
So we go on with the second option which was to tilt the beam with SIB1_M3 mirror. We managed to shift PR by about 1 mm but shifting the PR tilted the green beam and we did not manage to inject it back into the fibers on EIB.
We made a step back which correspond to a fianl shift of the PR of 0.4 mm, and manage to find the green back. The shift on the PZT are quite large, and we had to move them both in vertical and horizontal (see attached picture). We were actualy quite close to the safety clipping of the PZT dynamic (+- 5V) and this might explain why it was impossible to inject the beam after the step of 1 mm.

The activitty will go on tomorrow morning, we may have to act manually on the mirrors on EIB to get back some dynamic for the PZT.
More detail will be given in a further entry.

Images attached to this report
Comments to this report:
capocasa, chiummo, gosselin, melo, ruggi, spinicelli - 2:38 Thursday 28 April 2022 (55640) Print this report

We continued the PR recentering during the morning shift. 

At first, we checked the dynamics of the controllers of the pzt used to recover the green transmission: we found that there is enough margin to control the green beam injection on EIB up to +-10V applied. 

We thus continued in the same direction with M3. After some work, we also recovered the green beam and checked the centering of the PR at the CITF.

Then, we operated with M4 in order to compensate the excess of shift of the beam induced by the M3 movement. Once we found the good direction, we iterated once again the movement of M3+M4, recovered the green (verifying that we do not exceed the maximum correction) and studied the centering at the CITF. 

At the end of the day, we got a total shift of the PR of 1mm (compared to the 4mm required from previous analysis). 

The whole movements of the pico actuator from the beginning are the following:

  • M3_TY backward 3900 steps
  • M4_TX forward 800 steps

The activity will continue tomorrow morning. A more detailed report will follow once the centering of the PR will be finalized. 



gosselin, spinicelli - 10:11 Thursday 28 April 2022 (55642) Print this report

About the actuation of the ALS PZT on EIB, we checked that adding 1 V on the setpoint on the vpm is acutally resulting in 5 V on the PZT driver (factor 10 more for the pzt driver, factor 2 less for the differential output).
The PZT driver has an offset of 50 V and a dynamic of +- 70 V around this point.
This means that we could actuate with +- 14 V. The safety clipping for the correction in ACL was +- 5V, we increased it to +- 10 V. 

bersanetti, capocasa, casanueva, derossi, gosselin, gouaty, mantovani, spinicelli - 18:36 Friday 29 April 2022 (55657) Print this report

Today we performed a third displacement of the input beam, corresponding to another 1mm of shift of the PR vertical position. 

Similarly to what we did in the last days, we applied 4000 steps backward on M3_Ty and 1200 steps forward on M4_Tx. Thanks to the re-alignment of yesterday morning, we could recover the green beam transmission by using pzt on EIB. 

We then checked the coupling of PR TX on PRCL compared to the previous step of yesterday (fig. 1) and to before the recentering (fig. 2). The coupling since the beginning seems to be consistent with a shift of ~ 3mm of the beam.

As a result of the recentering, both sideband increase by ~10% (see fig. 3).

To be noted that, as a side effect of the PR recentering, the RFC_PD1 resulted partially misaligned. R. Gouaty recovered a good alignment of the photodiode acting on the picomotors on SIB2.  

Images attached to this comment
Search Help